Help please:
Please show and explain work
Question 23 (Worth 2 points)
Find the product (x+5)(x+7).
Question 29 (Worth 2 points)
Solve the inequality and describe the graph of the solution.
-3x + 4 > 7
Question 31 (Worth 2 points)
Which of the following relations is NOT a function?
(2,1), (1,-2), (-3,2), (2,-3)
(2,-1), (-1,2), (-2,3), (3,-2)
(-3,-2), (-1,0), (0,1), (1,2)

(-2,1), (-1,2), (1,1), (2,3)
Question 36 (Worth 2.0 points)
What is the degree of this polynomial:
3x3 - 2y4 + 2xy2
Question 41 (Worth 2 points)
Solve the following expression when x = 4, and y = -3.
2(x-2y)
Answer blank 1:
Answer not accepted. Expected Answer: 20
Question Information: 2[4-2(-3)]=
2(4+6)=
2(10)=
20
Question 44 (Worth 2 points)
Find: - 8xy(4x2y + 7x - 6y2)
Question 46 (Worth 2 points)
Solve the following system of equations using the substitution method:
4x - y = -6
x + 2y = 3


Answer by checkley75(3666)

(X+5)(X+7)=X^2+5X+7X+35=X^2+12X+35
3X+4>7, 3X>7-4, X>3/3, X>1 (NO GRAPH)
THE REST I LEAVE UP TO YOU.
WE SOLVE PROBLEMS. WE DON'T DO HOMEWORK!